You are trying to decide a present worth of a contract. You will receive $10,000 when the contract is signed, a $20,000 payme only numbers 2,3, and 6
Answer #1

2.

Quarterly rate = 4%/4 = 1%

Let, number of quarters required = n

Then,

105000 = 15000 + 3005*(1-1/(1+1%)^n)/1%

Or,

(1-1/1.01^n) = 1%*((105000-15000)/3005)

(1-1/1.01^n) = .2995

1/1.01^n = 1-.2995

1/(1-.2995) = 1.01^n

1.01^n = 1.42755

n = log(1.42755)/log (1.01)

n = 35.78 quarters

So, number of years required = 8.95 years or 9 years

================

3.

Number of semiannual payments left = 36 semiannual payment (18 years)

Semiannual coupon payment = (4%/2)*1000 = $20

Semiannual interest rate = 6.09%/2 = 3.045%

So,

Price of the bond = 20*(1-1/(1+3.045%)^36)/3.045% + 1000/1.03045^36

Price of the bond = $773.38 or $773

So, price of the bond should be $773.38 or $773.

====

Pl. repost other unanswered questions for their proper answers!